A car leaves a town traveling at 72 km/h. An hour later, a second car leaves the same town, on the same road traveling at 90 kn/h. In how many hours will the second car overtake the first?: A car leaves a town traveling at 72 km/h. An hour later, a second car leaves the same town, on the same road traveling at 90 kn/h. In how many hours will the second car overtake the first?
Answer by josmiceli(2170)

They both will have gone the same distance
when they meet.
Let d= the distance they both travel
Let t= the time elapsed during the 1st car's trip
For 1st car:
d = r*t
d = 72t
For the 2nd car
d = 90*(t - 1)
Since both d's are equal,
72t = 90*(t - 1)
72t = 90t - 90
18t = 90
t = 5hrs
The 2nd car's elapsed time is
t - 1
5 - 1 = 4hrs
The 2nd car overtakes the 1st car in 4 hours
Check the answer:
For the 1st car
d = 72*5
d = 360km
For the 2nd car
d = 90*4
d= 360km
OK